OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2 x + y = –8 3x + y = –2 x + y = –8 5x + 3y = 14 7x + 7y = 28 3x + y = –2 7x + 7y = 28 5x + 3y = 10

OpenStudy (mathstudent55):

To eliminate z using the first and second equations, multiply the second equation by 3 and add to the first one. To eliminate z using the second and third equations, just add them together.

OpenStudy (anonymous):

x + y - 3z = - 8 2x + 2y +z = 12 3x + y - z = - 2 ---------------- x + y - 3z = - 8 2x + 2y +z = 12 ---> (3)2x + 2y + z = 12 ----------------- x + y - 3z = - 8 6x + 6y + 3z = 36 (result of multiplying by 3) ----------------- 7x + 7y = 10 (result from equation 1 and 2) 2x + 2y + z = 12 3x + y - z = - 2 ---------------- 5x + 3y = 10 (result from equation 2 and 3) ANSWER : 7x + 7y = 10 ...............5x + 3y = 10

hero (hero):

That isn't even one of the answer choices

OpenStudy (mathstudent55):

you mean 7x + 7y = 28 and 5x + 3y = 10 which is D

OpenStudy (anonymous):

my bad...it is D. Sorry for messing up.